OpenStudy (jessiegonzales):

MEDAL!! Write an equation of the line, in point-slope form, that passes through the two given points. (-2,15) (9,-18) y-2=-3(x-15) y-2=-1/3(x+15) y-15=-3(x+2) y-15=-1/3(x+2)

OpenStudy (anonymous):

a) find the slope what do you get?

OpenStudy (jessiegonzales):

ok so 15-18 -2-9

OpenStudy (anonymous):

oh no

OpenStudy (jessiegonzales):

-3/-11

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

\[\huge \frac{\color{red}{y_2}-\color{green}{y_1}}{(\color{blue}{x_2}-\color{orange}{x_1})}\]

OpenStudy (anonymous):

\[\color{orange}{x_1}=\color{orange}{9}\] \[\color{blue}{x_2}=\color{blue}{-2}\] \[\color{green}{y_1}=\color{green}{15}\] \[\color{red}{y_2}=\color{red}{-18}\]

OpenStudy (anonymous):

remember the sign comes with the number

OpenStudy (jessiegonzales):

7/-3

OpenStudy (anonymous):

so for example your numerator should be \[15-(-18)\] not \[15-18\]

OpenStudy (jessiegonzales):

oh

OpenStudy (anonymous):

and similarly, your denominator should be \[-2-9\]

OpenStudy (jessiegonzales):

which is -7

OpenStudy (anonymous):

\[-2-9\neq -7\]

OpenStudy (anonymous):

\[2-9=-7\] but not \(-2-9\)

OpenStudy (jessiegonzales):

oh I'm confused now

OpenStudy (anonymous):

i think the arithmetic is throwing you off lets do the top first what is \[15-(-18)\]?

OpenStudy (anonymous):

if that is not clear, let me know and i will tell you

OpenStudy (jessiegonzales):

33

OpenStudy (anonymous):

yes ok good now on to the denominator

OpenStudy (anonymous):

denominator is \[-2-9\]

OpenStudy (jessiegonzales):

ok

OpenStudy (anonymous):

do you know what that is hint: it is NOT \(-7\)

OpenStudy (jessiegonzales):

-11

OpenStudy (jessiegonzales):

33/-11

OpenStudy (anonymous):

whew got it

OpenStudy (anonymous):

right, and \[\frac{33}{-11}=?\]

OpenStudy (jessiegonzales):

-1/3

OpenStudy (anonymous):

hmmm no

OpenStudy (jessiegonzales):

sorry -3

OpenStudy (anonymous):

right

OpenStudy (jessiegonzales):

a is the answer

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] you have \(m=-3\)

OpenStudy (anonymous):

they are using the first point \((-2,15)\) for \((x_1,y_1)\)

OpenStudy (jessiegonzales):

c is the answer thankyou so much

OpenStudy (anonymous):

yw